# The N3LO Twist-2 Matching of Helicity TMDs and SIDIS $q_{\ast}$ Spectrum

### v3

@@ 4.4.2 Time-like results in $\overline{\text{MS}}$}
$\Delta P_{gq}^{T,\text{s}(2)}=C_{F}^{3}\bigg[\left(-\frac{48}{z}\right)H_{-1,0}+447H_{0}+626H_{1}+32H_{-1}\zeta_{2}+24H_{0}\zeta_{2}-112H_{1}\zeta_{2}-424\zeta_{2}-400H_{0}\zeta_{3}+64H_{1}\zeta_{3}-224\zeta_{3}+84\zeta_{4}+48H_{-1,0}+192\zeta_{2}H_{0,-1}-128\zeta_{2}H_{0,0}+174H_{0,0}+400\zeta_{2}H_{0,1}+40H_{0,1}+128\zeta_{2}H_{1,0}+304H_{1,0}+96\zeta_{2}H_{1,1}+240H_{1,1}+64H_{-1,-1,0}+64H_{-1,0,0}-168H_{0,0,0}-344H_{0,0,1}+40H_{0,1,0}-40H_{0,1,1}+80H_{1,0,0}-16H_{1,0,1}+80H_{1,1,0}+96H_{1,1,1}-128H_{-1,-1,0,0}-64H_{-1,0,-1,0}+32H_{-1,0,0,0}+384H_{0,-1,-1,0}-128H_{0,-1,0,0}-192H_{0,0,-1,0}+192H_{0,0,0,0}+128H_{0,0,0,1}+176H_{0,0,1,0}+240H_{0,0,1,1}+80H_{0,1,0,0}-144H_{0,1,0,1}-80H_{0,1,1,0}-208H_{0,1,1,1}-64H_{1,0,-1,0}-192H_{1,0,0,0}-256H_{1,0,0,1}+32H_{1,0,1,0}-128H_{1,0,1,1}-192H_{1,1,0,0}-96H_{1,1,0,1}+32H_{1,1,1,0}+z\left(-60\zeta_{2}H_{0}-40\zeta_{3}H_{0}-221H_{0}-592H_{1}+32H_{-1}\zeta_{2}+76H_{1}\zeta_{2}+44\zeta_{2}-32H_{1}\zeta_{3}+76\zeta_{3}-194\zeta_{4}+64H_{-1,0}+64\zeta_{2}H_{0,-1}-32\zeta_{2}H_{0,0}-73H_{0,0}-232\zeta_{2}H_{0,1}+244H_{0,1}-64\zeta_{2}H_{1,0}-248H_{1,0}-48\zeta_{2}H_{1,1}-184H_{1,1}+64H_{-1,-1,0}+64H_{-1,0,0}-144H_{0,-1,0}+108H_{0,0,0}+140H_{0,0,1}+56H_{0,1,0}+112H_{0,1,1}+28H_{1,0,0}+52H_{1,0,1}-44H_{1,1,0}-84H_{1,1,1}-64H_{-1,-1,0,0}-32H_{-1,0,-1,0}+16H_{-1,0,0,0}+128H_{0,-1,-1,0}-32H_{0,-1,0,0}-32H_{0,0,-1,0}-16H_{0,0,0,0}-64H_{0,0,0,1}-88H_{0,0,1,0}-120H_{0,0,1,1}-8H_{0,1,0,0}+72H_{0,1,0,1}+40H_{0,1,1,0}+104H_{0,1,1,1}+32H_{1,0,-1,0}+96H_{1,0,0,0}+128H_{1,0,0,1}-16H_{1,0,1,0}+64H_{1,0,1,1}+96H_{1,1,0,0}+48H_{1,1,0,1}-16H_{1,1,1,0}-32H_{1,1,1,1}-\frac{3133}{4}\right)+64H_{1,1,1,1}+\frac{1569}{2}\bigg]+N_{f}C_{F}^{2}\bigg[\left(-\frac{80}{3}\right)H_{1,0}+\left(-\frac{80}{3}\right)H_{1,0,0}+\left(-\frac{112}{9}\right)H_{1,1}+\left(-\frac{32}{3}\right)H_{0,1,1}+\left(-\frac{16}{3}\right)H_{0,0,0}+\left(-\frac{16}{3}\right)H_{1,0,1}+\left(-\frac{16}{3}\right)H_{1,1,0}+\frac{76}{27}H_{1}+\frac{16}{3}H_{1,1,1}+\frac{64}{3}H_{0,0,1}+\frac{224}{9}H_{0,1}+\frac{380}{3}H_{0,0}+\frac{4334}{27}H_{0}-16H_{1}\zeta_{2}-64H_{0,0,0,0}+z\left(\left(-\frac{328}{3}\right)H_{0,0,0}+\left(-\frac{2038}{27}\right)H_{0}+\left(-\frac{32}{3}\right)H_{0,0,1}+\left(-\frac{64}{9}\right)H_{0,1}+\left(-\frac{8}{3}\right)H_{1,1,1}+\frac{8}{3}H_{1,0,1}+\frac{8}{3}H_{1,1,0}+\frac{32}{9}H_{1,1}+\frac{112}{27}H_{1}+\frac{16}{3}H_{0,1,1}+\frac{40}{3}H_{1,0}+\frac{40}{3}H_{1,0,0}+8H_{1}\zeta_{2}+154H_{0,0}+32H_{0,0,0,0}-\frac{1567}{18}\right)+\frac{767}{9}\bigg]+C_{A}^{2}C_{F}\bigg[\left(-\frac{7520}{9}\right)\zeta_{2}+\left(-\frac{7376}{9}\right)H_{1,0}+\left(-\frac{1384}{3}\right)H_{-1,0,0}+\left(-\frac{904}{3}\right)H_{0,-1,0}+\left(-\frac{2248}{9}\right)H_{1,1}+\left(-\frac{2216}{9}\right)H_{-1,0}+\left(-\frac{268}{3}\right)H_{1,0,0}+\frac{8}{3}H_{0,1,1}+\frac{88}{3}H_{1,1,1}+\frac{112}{3}H_{1,1,0}+\frac{932}{3}H_{0,0,1}+\frac{5584}{9}H_{0,0}+\frac{2216}{3}H_{0,0,0}+\frac{7624}{9}H_{0,1}+\frac{28256}{27}H_{0}+\frac{32362}{27}H_{1}+\left(-\frac{24}{z}\right)H_{-1,0}+\left(-\frac{1468\zeta_{2}}{3}\right)H_{0}+104H_{-1}\zeta_{2}+264H_{1}\zeta_{2}-16H_{-1}\zeta_{3}+480H_{0}\zeta_{3}+656H_{1}\zeta_{3}-712\zeta_{3}+1570\zeta_{4}-32\zeta_{2}H_{-1,-1}$ (134)
$160\zeta_{2}H_{-1,0}-16\zeta_{2}H_{0,-1}-312\zeta_{2}H_{0,0}+496\zeta_{2}H_{0,1}-128\zeta_{2}H_{1,0}+96\zeta_{2}H_{1,1}+336H_{-1,-1,0}+64H_{-1,0,1}+288H_{0,1,0}-96H_{1,0,1}-192H_{-1,-1,-1,0}+192H_{-1,-1,0,0}-64H_{-1,-1,0,1}+352H_{-1,0,-1,0}-336H_{-1,0,0,0}+64H_{-1,0,0,1}-192H_{-1,0,1,0}-64H_{-1,0,1,1}+96H_{0,-1,-1,0}-272H_{0,-1,0,0}+64H_{0,-1,0,1}-496H_{0,0,-1,0}+1264H_{0,0,0,0}-184H_{0,0,0,1}+960H_{0,0,1,0}+320H_{0,0,1,1}-152H_{0,1,0,0}-352H_{0,1,0,1}-224H_{0,1,1,0}-160H_{0,1,1,1}-96H_{1,0,-1,0}+560H_{1,0,0,0}+160H_{1,0,0,1}+640H_{1,0,1,0}+64H_{1,0,1,1}+160H_{1,1,0,0}+128H_{1,1,0,1}+192H_{1,1,1,0}+z\left(\left(-\frac{31412}{27}\right)H_{1}+\left(-\frac{20644}{27}\right)H_{0}+\left(-\frac{1472}{3}\right)H_{-1,0,0}+\left(-\frac{992}{3}\right)H_{0,-1,0}+\left(-\frac{2752}{9}\right)H_{-1,0}+\left(-\frac{580}{3}\right)H_{0,0,1}+\left(-\frac{484}{3}\right)H_{0,1,1}+\left(-\frac{328}{3}\right)H_{0,0,0}+\left(-\frac{200}{3}\right)H_{1,1,0}+\left(-\frac{172}{3}\right)H_{1,0,0}+\left(-\frac{44}{3}\right)H_{1,1,1}+\frac{1712}{9}H_{1,1}+\frac{5380}{9}H_{0,1}+\frac{5944}{9}H_{1,0}+\frac{10172}{9}H_{0,0}+104H_{-1}\zeta_{2}-156H_{0}\zeta_{2}-352H_{1}\zeta_{2}-796\zeta_{2}-8H_{-1}\zeta_{3}+176H_{0}\zeta_{3}-328H_{1}\zeta_{3}-464\zeta_{3}+143\zeta_{4}-16\zeta_{2}H_{-1,-1}+80\zeta_{2}H_{-1,0}-72\zeta_{2}H_{0,-1}+76\zeta_{2}H_{0,0}+24\zeta_{2}H_{0,1}+64\zeta_{2}H_{1,0}-48\zeta_{2}H_{1,1}+336H_{-1,-1,0}+64H_{-1,0,1}-848H_{0,1,0}+96H_{1,0,1}-96H_{-1,-1,-1,0}+96H_{-1,-1,0,0}-32H_{-1,-1,0,1}+176H_{-1,0,-1,0}-168H_{-1,0,0,0}+32H_{-1,0,0,1}-96H_{-1,0,1,0}-32H_{-1,0,1,1}-80H_{0,-1,-1,0}+120H_{0,-1,0,0}+32H_{0,-1,0,1}+56H_{0,0,-1,0}-64H_{0,0,0,0}-76H_{0,0,0,1}+192H_{0,0,1,0}+64H_{0,0,1,1}+116H_{0,1,0,0}+32H_{0,1,0,1}+64H_{0,1,1,0}+32H_{0,1,1,1}+48H_{1,0,-1,0}-280H_{1,0,0,0}-80H_{1,0,0,1}-320H_{1,0,1,0}-32H_{1,0,1,1}-80H_{1,1,0,0}-64H_{1,1,0,1}-96H_{1,1,1,0}-32H_{1,1,1,1}-\frac{13346}{9}\right)+64H_{1,1,1,1}+\frac{13840}{9}\bigg]+C_{A}C_{F}^{2}\bigg[\left(-\frac{52072}{27}\right)H_{1}+\left(-\frac{10868}{9}\right)H_{0,1}+\left(-\frac{1466}{3}\right)H_{0,0}+\left(-\frac{1136}{3}\right)H_{0,0,0}+\left(-\frac{3719}{27}\right)H_{0}+\left(-\frac{376}{3}\right)H_{1,1,1}+\left(-\frac{368}{3}\right)H_{1,1,0}+\left(-\frac{184}{3}\right)H_{0,0,1}+\left(-\frac{344}{9}\right)H_{1,1}+\left(-\frac{88}{3}\right)H_{1,0}+\frac{56}{3}H_{0,1,1}+\frac{272}{3}H_{1,0,0}+\frac{352}{3}H_{1,0,1}+\frac{1}{z}\left(-72H_{-1}\zeta_{2}-24H_{1}\zeta_{2}+72H_{-1,0}-48H_{-1,-1,0}+48H_{-1,0,0}+48H_{-1,0,1}\right)-256H_{-1}\zeta_{2}+296H_{0}\zeta_{2}-240H_{1}\zeta_{2}+716\zeta_{2}-176H_{-1}\zeta_{3}-752H_{0}\zeta_{3}-816H_{1}\zeta_{3}+728\zeta_{3}-2144\zeta_{4}+288\zeta_{2}H_{-1,-1}-192\zeta_{2}H_{-1,0}+136H_{-1,0}-32\zeta_{2}H_{0,-1}-80\zeta_{2}H_{0,0}-912\zeta_{2}H_{0,1}+32\zeta_{2}H_{1,0}-192\zeta_{2}H_{1,1}-256H_{-1,-1,0}+144H_{-1,0,0}+128H_{-1,0,1}+96H_{0,-1,0}-472H_{0,1,0}+192H_{-1,-1,-1,0}+64H_{-1,-1,0,0}-192H_{-1,-1,0,1}-160H_{-1,0,-1,0}-16H_{-1,0,0,0}-64H_{-1,0,0,1}+128H_{-1,0,1,0}+64H_{-1,0,1,1}-192H_{0,-1,-1,0}-64H_{0,-1,0,0}-64H_{0,-1,0,1}+96H_{0,0,-1,0}+128H_{0,0,0,0}+752H_{0,0,0,1}-688H_{0,0,1,0}-432H_{0,0,1,1}+416H_{0,1,0,0}+624H_{0,1,0,1}+432H_{0,1,1,0}+368H_{0,1,1,1}+96H_{1,0,-1,0}-48H_{1,0,0,0}+224H_{1,0,0,1}-544H_{1,0,1,0}+64H_{1,0,1,1}+160H_{1,1,0,0}-32H_{1,1,0,1}-224H_{1,1,1,0}-128H_{1,1,1,1}+z\left(\left(-\frac{6224}{9}\right)H_{0,1}+\left(-\frac{688}{3}\right)H_{0,0,1}+\left(-\frac{452}{3}\right)H_{1,0,1}+\left(-\frac{448}{3}\right)H_{1,0,0}+\frac{56}{3}H_{1,0}+\frac{68}{3}H_{0,1,1}+\frac{376}{9}H_{1,1}+\frac{296}{3}H_{1,1,1}+\frac{340}{3}H_{1,1,0}+\frac{364}{3}H_{0,0,0}+\frac{50168}{27}H_{1}+\frac{57247}{27}H_{0}-184H_{-1}\zeta_{2}+72H_{0}\zeta_{2}+380H_{1}\zeta_{2}+1176\zeta_{2}-88H_{-1}\zeta_{3}-136H_{0}\zeta_{3}+408H_{1}\zeta_{3}+368\zeta_{3}-152\zeta_{4}+144\zeta_{2}H_{-1,-1}-96\zeta_{2}H_{-1,0}+64H_{-1,0}-64\zeta_{2}H_{0,-1}-8\zeta_{2}H_{0,0}-889H_{0,0}+136\zeta_{2}H_{0,1}-16\zeta_{2}H_{1,0}+96\zeta_{2}H_{1,1}-208H_{-1,-1,0}+24H_{-1,0,0}+80H_{-1,0,1}+264H_{0,-1,0}+528H_{0,1,0}+96H_{-1,-1,-1,0}+32H_{-1,-1,0,0}-96H_{-1,-1,0,1}-80H_{-1,0,-1,0}-8H_{-1,0,0,0}-32H_{-1,0,0,1}+64H_{-1,0,1,0}+32H_{-1,0,1,1}-64H_{0,-1,-1,0}-16H_{0,-1,0,0}+32H_{0,-1,0,1}-16H_{0,0,-1,0}+32H_{0,0,0,0}+168H_{0,0,0,1}-104H_{0,0,1,0}-8H_{0,0,1,1}-160H_{0,1,0,0}-168H_{0,1,0,1}-168H_{0,1,1,0}-136H_{0,1,1,1}-48H_{1,0,-1,0}+24H_{1,0,0,0}-112H_{1,0,0,1}+272H_{1,0,1,0}-32H_{1,0,1,1}-80H_{1,1,0,0}+16H_{1,1,0,1}+112H_{1,1,1,0}+64H_{1,1,1,1}-\frac{2737}{36}\right)+\frac{1577}{18}\bigg]$ (135)
$N_{f}C_{A}C_{F}\bigg[\left(-\frac{728}{27}\right)H_{0}+\left(-\frac{112}{9}\right)H_{-1,0}+\left(-\frac{32}{3}\right)H_{-1,0,0}+\left(-\frac{32}{3}\right)H_{0,-1,0}+\left(-\frac{32}{3}\right)H_{0,0,1}+\left(-\frac{88}{9}\right)H_{0,1}+\left(-\frac{16}{3}\right)H_{1,1,1}+\frac{8}{9}\zeta_{2}+\frac{32}{27}H_{1}+\frac{16}{3}H_{0,1,1}+\frac{32}{3}H_{1,1,0}+\frac{112}{9}H_{1,1}+\frac{368}{9}H_{1,0}+\frac{160}{3}H_{1,0,0}+\frac{584}{9}H_{0,0}+\frac{256}{3}H_{0,0,0}+\left(-\frac{32\zeta_{2}}{3}\right)H_{0}+32H_{1}\zeta_{2}-80\zeta_{3}+z\left(\left(-\frac{356}{9}\right)H_{0,0}+\left(-\frac{80}{3}\right)H_{1,0,0}+\left(-\frac{208}{9}\right)H_{1,0}+\left(-\frac{518}{27}\right)H_{0}+\left(-\frac{184}{27}\right)H_{1}+\left(-\frac{16}{3}\right)H_{-1,0,0}+\left(-\frac{16}{3}\right)H_{0,-1,0}+\left(-\frac{16}{3}\right)H_{1,1,0}+\left(-\frac{32}{9}\right)H_{-1,0}+\left(-\frac{32}{9}\right)H_{1,1}+\left(-\frac{8}{3}\right)H_{0,1,1}+\frac{8}{3}H_{1,1,1}+\frac{32}{9}H_{0,1}+\frac{16}{3}H_{0,0,0}+\frac{16}{3}H_{0,0,1}-16H_{1}\zeta_{2}-16\zeta_{2}+32\zeta_{3}+\frac{416}{9}\right)-\frac{526}{9}\bigg]\,$, (136)

### v4

@@ 4.4.2 Time-like results in $\overline{\text{MS}}$}
$\Delta P_{gq}^{T,\text{s}(2)}=C_{F}^{3}\bigg[\left(-\frac{48}{z}\right)H_{-1,0}+447H_{0}+626H_{1}+32H_{-1}\zeta_{2}+24H_{0}\zeta_{2}-112H_{1}\zeta_{2}-424\zeta_{2}-400H_{0}\zeta_{3}+64H_{1}\zeta_{3}-224\zeta_{3}+84\zeta_{4}+48H_{-1,0}+192\zeta_{2}H_{0,-1}-128\zeta_{2}H_{0,0}+174H_{0,0}+400\zeta_{2}H_{0,1}+40H_{0,1}+128\zeta_{2}H_{1,0}+304H_{1,0}+96\zeta_{2}H_{1,1}+240H_{1,1}+64H_{-1,-1,0}+64H_{-1,0,0}-168H_{0,0,0}-344H_{0,0,1}+40H_{0,1,0}-40H_{0,1,1}+80H_{1,0,0}-16H_{1,0,1}+80H_{1,1,0}+96H_{1,1,1}-128H_{-1,-1,0,0}-64H_{-1,0,-1,0}+32H_{-1,0,0,0}+384H_{0,-1,-1,0}-128H_{0,-1,0,0}-192H_{0,0,-1,0}+192H_{0,0,0,0}+128H_{0,0,0,1}+176H_{0,0,1,0}+240H_{0,0,1,1}+80H_{0,1,0,0}-144H_{0,1,0,1}-80H_{0,1,1,0}-208H_{0,1,1,1}-64H_{1,0,-1,0}-192H_{1,0,0,0}-256H_{1,0,0,1}+32H_{1,0,1,0}-128H_{1,0,1,1}-192H_{1,1,0,0}-96H_{1,1,0,1}+32H_{1,1,1,0}+z\left(-60\zeta_{2}H_{0}-40\zeta_{3}H_{0}-221H_{0}-592H_{1}+32H_{-1}\zeta_{2}+76H_{1}\zeta_{2}+44\zeta_{2}-32H_{1}\zeta_{3}+76\zeta_{3}-194\zeta_{4}+64H_{-1,0}+64\zeta_{2}H_{0,-1}-32\zeta_{2}H_{0,0}-73H_{0,0}-232\zeta_{2}H_{0,1}+244H_{0,1}-64\zeta_{2}H_{1,0}-248H_{1,0}-48\zeta_{2}H_{1,1}-184H_{1,1}+64H_{-1,-1,0}+64H_{-1,0,0}-144H_{0,-1,0}+108H_{0,0,0}+140H_{0,0,1}+56H_{0,1,0}+112H_{0,1,1}+28H_{1,0,0}+52H_{1,0,1}-44H_{1,1,0}-84H_{1,1,1}-64H_{-1,-1,0,0}-32H_{-1,0,-1,0}+16H_{-1,0,0,0}+128H_{0,-1,-1,0}-32H_{0,-1,0,0}-32H_{0,0,-1,0}-16H_{0,0,0,0}-64H_{0,0,0,1}-88H_{0,0,1,0}-120H_{0,0,1,1}-8H_{0,1,0,0}+72H_{0,1,0,1}+40H_{0,1,1,0}+104H_{0,1,1,1}+32H_{1,0,-1,0}+96H_{1,0,0,0}+128H_{1,0,0,1}-16H_{1,0,1,0}+64H_{1,0,1,1}+96H_{1,1,0,0}+48H_{1,1,0,1}-16H_{1,1,1,0}-32H_{1,1,1,1}-\frac{3133}{4}\right)+64H_{1,1,1,1}+\frac{1569}{2}\bigg]+N_{f}C_{F}^{2}\bigg[\left(-\frac{80}{3}\right)H_{1,0}+\left(-\frac{80}{3}\right)H_{1,0,0}+\left(-\frac{112}{9}\right)H_{1,1}+\left(-\frac{32}{3}\right)H_{0,1,1}+\left(-\frac{16}{3}\right)H_{0,0,0}+\left(-\frac{16}{3}\right)H_{1,0,1}+\left(-\frac{16}{3}\right)H_{1,1,0}+\frac{76}{27}H_{1}+\frac{16}{3}H_{1,1,1}+\frac{64}{3}H_{0,0,1}+\frac{224}{9}H_{0,1}+\frac{380}{3}H_{0,0}+\frac{4334}{27}H_{0}-16H_{1}\zeta_{2}-64H_{0,0,0,0}+z\left(\left(-\frac{328}{3}\right)H_{0,0,0}+\left(-\frac{2038}{27}\right)H_{0}+\left(-\frac{32}{3}\right)H_{0,0,1}+\left(-\frac{64}{9}\right)H_{0,1}+\left(-\frac{8}{3}\right)H_{1,1,1}+\frac{8}{3}H_{1,0,1}+\frac{8}{3}H_{1,1,0}+\frac{32}{9}H_{1,1}+\frac{112}{27}H_{1}+\frac{16}{3}H_{0,1,1}+\frac{40}{3}H_{1,0}+\frac{40}{3}H_{1,0,0}+8H_{1}\zeta_{2}+154H_{0,0}+32H_{0,0,0,0}-\frac{1567}{18}\right)+\frac{767}{9}\bigg]+C_{A}^{2}C_{F}\bigg[\left(-\frac{7520}{9}\right)\zeta_{2}+\left(-\frac{7376}{9}\right)H_{1,0}+\left(-\frac{1384}{3}\right)H_{-1,0,0}+\left(-\frac{904}{3}\right)H_{0,-1,0}+\left(-\frac{2248}{9}\right)H_{1,1}+\left(-\frac{2216}{9}\right)H_{-1,0}+\left(-\frac{268}{3}\right)H_{1,0,0}+\frac{8}{3}H_{0,1,1}+\frac{88}{3}H_{1,1,1}+\frac{112}{3}H_{1,1,0}+\frac{932}{3}H_{0,0,1}+\frac{5584}{9}H_{0,0}+\frac{2216}{3}H_{0,0,0}+\frac{7624}{9}H_{0,1}+\frac{28256}{27}H_{0}+\frac{32362}{27}H_{1}+\left(-\frac{24}{z}\right)H_{-1,0}+\left(-\frac{1468\zeta_{2}}{3}\right)H_{0}+104H_{-1}\zeta_{2}+264H_{1}\zeta_{2}-16H_{-1}\zeta_{3}+480H_{0}\zeta_{3}+656H_{1}\zeta_{3}-712\zeta_{3}+1570\zeta_{4}-32\zeta_{2}H_{-1,-1}$
$160\zeta_{2}H_{-1,0}-16\zeta_{2}H_{0,-1}-312\zeta_{2}H_{0,0}+496\zeta_{2}H_{0,1}-128\zeta_{2}H_{1,0}+96\zeta_{2}H_{1,1}+336H_{-1,-1,0}+64H_{-1,0,1}+288H_{0,1,0}-96H_{1,0,1}-192H_{-1,-1,-1,0}+192H_{-1,-1,0,0}-64H_{-1,-1,0,1}+352H_{-1,0,-1,0}-336H_{-1,0,0,0}+64H_{-1,0,0,1}-192H_{-1,0,1,0}-64H_{-1,0,1,1}+96H_{0,-1,-1,0}-272H_{0,-1,0,0}+64H_{0,-1,0,1}-496H_{0,0,-1,0}+1264H_{0,0,0,0}-184H_{0,0,0,1}+960H_{0,0,1,0}+320H_{0,0,1,1}-152H_{0,1,0,0}-352H_{0,1,0,1}-224H_{0,1,1,0}-160H_{0,1,1,1}-96H_{1,0,-1,0}+560H_{1,0,0,0}+160H_{1,0,0,1}+640H_{1,0,1,0}+64H_{1,0,1,1}+160H_{1,1,0,0}+128H_{1,1,0,1}+192H_{1,1,1,0}+z\left(\left(-\frac{31412}{27}\right)H_{1}+\left(-\frac{20644}{27}\right)H_{0}+\left(-\frac{1472}{3}\right)H_{-1,0,0}+\left(-\frac{992}{3}\right)H_{0,-1,0}+\left(-\frac{2752}{9}\right)H_{-1,0}+\left(-\frac{580}{3}\right)H_{0,0,1}+\left(-\frac{484}{3}\right)H_{0,1,1}+\left(-\frac{328}{3}\right)H_{0,0,0}+\left(-\frac{200}{3}\right)H_{1,1,0}+\left(-\frac{172}{3}\right)H_{1,0,0}+\left(-\frac{44}{3}\right)H_{1,1,1}+\frac{1712}{9}H_{1,1}+\frac{5380}{9}H_{0,1}+\frac{5944}{9}H_{1,0}+\frac{10172}{9}H_{0,0}+104H_{-1}\zeta_{2}-156H_{0}\zeta_{2}-352H_{1}\zeta_{2}-796\zeta_{2}-8H_{-1}\zeta_{3}+176H_{0}\zeta_{3}-328H_{1}\zeta_{3}-464\zeta_{3}+143\zeta_{4}-16\zeta_{2}H_{-1,-1}+80\zeta_{2}H_{-1,0}-72\zeta_{2}H_{0,-1}+76\zeta_{2}H_{0,0}+24\zeta_{2}H_{0,1}+64\zeta_{2}H_{1,0}-48\zeta_{2}H_{1,1}+336H_{-1,-1,0}+64H_{-1,0,1}-848H_{0,1,0}+96H_{1,0,1}-96H_{-1,-1,-1,0}+96H_{-1,-1,0,0}-32H_{-1,-1,0,1}+176H_{-1,0,-1,0}-168H_{-1,0,0,0}+32H_{-1,0,0,1}-96H_{-1,0,1,0}-32H_{-1,0,1,1}-80H_{0,-1,-1,0}+120H_{0,-1,0,0}+32H_{0,-1,0,1}+56H_{0,0,-1,0}-64H_{0,0,0,0}-76H_{0,0,0,1}+192H_{0,0,1,0}+64H_{0,0,1,1}+116H_{0,1,0,0}+32H_{0,1,0,1}+64H_{0,1,1,0}+32H_{0,1,1,1}+48H_{1,0,-1,0}-280H_{1,0,0,0}-80H_{1,0,0,1}-320H_{1,0,1,0}-32H_{1,0,1,1}-80H_{1,1,0,0}-64H_{1,1,0,1}-96H_{1,1,1,0}-32H_{1,1,1,1}-\frac{13346}{9}\right)+64H_{1,1,1,1}+\frac{13840}{9}\bigg]+C_{A}C_{F}^{2}\bigg[\left(-\frac{52072}{27}\right)H_{1}+\left(-\frac{10868}{9}\right)H_{0,1}+\left(-\frac{1466}{3}\right)H_{0,0}+\left(-\frac{1136}{3}\right)H_{0,0,0}+\left(-\frac{3719}{27}\right)H_{0}+\left(-\frac{376}{3}\right)H_{1,1,1}+\left(-\frac{368}{3}\right)H_{1,1,0}+\left(-\frac{184}{3}\right)H_{0,0,1}+\left(-\frac{344}{9}\right)H_{1,1}+\left(-\frac{88}{3}\right)H_{1,0}+\frac{56}{3}H_{0,1,1}+\frac{272}{3}H_{1,0,0}+\frac{352}{3}H_{1,0,1}+\frac{1}{z}\left(-72H_{-1}\zeta_{2}-24H_{1}\zeta_{2}+72H_{-1,0}-48H_{-1,-1,0}+48H_{-1,0,0}+48H_{-1,0,1}\right)-256H_{-1}\zeta_{2}+296H_{0}\zeta_{2}-240H_{1}\zeta_{2}+716\zeta_{2}-176H_{-1}\zeta_{3}-752H_{0}\zeta_{3}-816H_{1}\zeta_{3}+728\zeta_{3}-2144\zeta_{4}+288\zeta_{2}H_{-1,-1}-192\zeta_{2}H_{-1,0}+136H_{-1,0}-32\zeta_{2}H_{0,-1}-80\zeta_{2}H_{0,0}-912\zeta_{2}H_{0,1}+32\zeta_{2}H_{1,0}-192\zeta_{2}H_{1,1}-256H_{-1,-1,0}+144H_{-1,0,0}+128H_{-1,0,1}+96H_{0,-1,0}-472H_{0,1,0}+192H_{-1,-1,-1,0}+64H_{-1,-1,0,0}-192H_{-1,-1,0,1}-160H_{-1,0,-1,0}-16H_{-1,0,0,0}-64H_{-1,0,0,1}+128H_{-1,0,1,0}+64H_{-1,0,1,1}-192H_{0,-1,-1,0}-64H_{0,-1,0,0}-64H_{0,-1,0,1}+96H_{0,0,-1,0}+128H_{0,0,0,0}+752H_{0,0,0,1}-688H_{0,0,1,0}-432H_{0,0,1,1}+416H_{0,1,0,0}+624H_{0,1,0,1}+432H_{0,1,1,0}+368H_{0,1,1,1}+96H_{1,0,-1,0}-48H_{1,0,0,0}+224H_{1,0,0,1}-544H_{1,0,1,0}+64H_{1,0,1,1}+160H_{1,1,0,0}-32H_{1,1,0,1}-224H_{1,1,1,0}-128H_{1,1,1,1}+z\left(\left(-\frac{6224}{9}\right)H_{0,1}+\left(-\frac{688}{3}\right)H_{0,0,1}+\left(-\frac{452}{3}\right)H_{1,0,1}+\left(-\frac{448}{3}\right)H_{1,0,0}+\frac{56}{3}H_{1,0}+\frac{68}{3}H_{0,1,1}+\frac{376}{9}H_{1,1}+\frac{296}{3}H_{1,1,1}+\frac{340}{3}H_{1,1,0}+\frac{364}{3}H_{0,0,0}+\frac{50168}{27}H_{1}+\frac{57247}{27}H_{0}-184H_{-1}\zeta_{2}+72H_{0}\zeta_{2}+380H_{1}\zeta_{2}+1176\zeta_{2}-88H_{-1}\zeta_{3}-136H_{0}\zeta_{3}+408H_{1}\zeta_{3}+368\zeta_{3}-152\zeta_{4}+144\zeta_{2}H_{-1,-1}-96\zeta_{2}H_{-1,0}+64H_{-1,0}-64\zeta_{2}H_{0,-1}-8\zeta_{2}H_{0,0}-889H_{0,0}+136\zeta_{2}H_{0,1}-16\zeta_{2}H_{1,0}+96\zeta_{2}H_{1,1}-208H_{-1,-1,0}+24H_{-1,0,0}+80H_{-1,0,1}+264H_{0,-1,0}+528H_{0,1,0}+96H_{-1,-1,-1,0}+32H_{-1,-1,0,0}-96H_{-1,-1,0,1}-80H_{-1,0,-1,0}-8H_{-1,0,0,0}-32H_{-1,0,0,1}+64H_{-1,0,1,0}+32H_{-1,0,1,1}-64H_{0,-1,-1,0}-16H_{0,-1,0,0}+32H_{0,-1,0,1}-16H_{0,0,-1,0}+32H_{0,0,0,0}+168H_{0,0,0,1}-104H_{0,0,1,0}-8H_{0,0,1,1}-160H_{0,1,0,0}-168H_{0,1,0,1}-168H_{0,1,1,0}-136H_{0,1,1,1}-48H_{1,0,-1,0}+24H_{1,0,0,0}-112H_{1,0,0,1}+272H_{1,0,1,0}-32H_{1,0,1,1}-80H_{1,1,0,0}+16H_{1,1,0,1}+112H_{1,1,1,0}+64H_{1,1,1,1}-\frac{2737}{36}\right)+\frac{1577}{18}\bigg]$
$N_{f}C_{A}C_{F}\bigg[\left(-\frac{728}{27}\right)H_{0}+\left(-\frac{112}{9}\right)H_{-1,0}+\left(-\frac{32}{3}\right)H_{-1,0,0}+\left(-\frac{32}{3}\right)H_{0,-1,0}+\left(-\frac{32}{3}\right)H_{0,0,1}+\left(-\frac{88}{9}\right)H_{0,1}+\left(-\frac{16}{3}\right)H_{1,1,1}+\frac{8}{9}\zeta_{2}+\frac{32}{27}H_{1}+\frac{16}{3}H_{0,1,1}+\frac{32}{3}H_{1,1,0}+\frac{112}{9}H_{1,1}+\frac{368}{9}H_{1,0}+\frac{160}{3}H_{1,0,0}+\frac{584}{9}H_{0,0}+\frac{256}{3}H_{0,0,0}+\left(-\frac{32\zeta_{2}}{3}\right)H_{0}+32H_{1}\zeta_{2}-80\zeta_{3}+z\left(\left(-\frac{356}{9}\right)H_{0,0}+\left(-\frac{80}{3}\right)H_{1,0,0}+\left(-\frac{208}{9}\right)H_{1,0}+\left(-\frac{518}{27}\right)H_{0}+\left(-\frac{184}{27}\right)H_{1}+\left(-\frac{16}{3}\right)H_{-1,0,0}+\left(-\frac{16}{3}\right)H_{0,-1,0}+\left(-\frac{16}{3}\right)H_{1,1,0}+\left(-\frac{32}{9}\right)H_{-1,0}+\left(-\frac{32}{9}\right)H_{1,1}+\left(-\frac{8}{3}\right)H_{0,1,1}+\frac{8}{3}H_{1,1,1}+\frac{32}{9}H_{0,1}+\frac{16}{3}H_{0,0,0}+\frac{16}{3}H_{0,0,1}-16H_{1}\zeta_{2}-16\zeta_{2}+32\zeta_{3}+\frac{416}{9}\right)-\frac{526}{9}\bigg]\,.$ (131)